OpenStudy (anonymous):

The sum of five times a number and four is equal to negative eleven. What is the number?

OpenStudy (cggurumanjunath):

5x+4 =-11

OpenStudy (cggurumanjunath):

5x = -11 -4 =-15 x= -15/5 =-3